xxcoffee:

y is it when i ask for help no one helps but when others ask they all get help ion get it

carmelle:

idk

dontsaymyname:

What question did you need help with? And maybe the question was tricky, or alot of work, or confusing???

xxcoffee:

no one gets on the post and i leave it up for ever

xxcoffee:

ding dong brother needs help with this

1 attachment
boredfr:

What type of help do you need?

xxcoffee:

but ion get it

jhonyy9:

23h - 12 = -10 like a first step add to both sides 12 so what will get ?

XioGonz:

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 23*h-12-(-10)=0

dontsaymyname:

Did my answers get deleted???

XioGonz:

Solve : 23h-2 = 0 Add 2 to both sides of the equation : 23h = 2 Divide both sides of the equation by 23: h= 2/23 = 0.087 And that rounds to 0.09

jhonyy9:

for the second one multiplie both sides by 17 to eliminate the denominator from left side so what will get ?

boredfr:

23h-12= -10 Step 1: Add 12 to both sides Step 2: Divide both sides by 23.

boredfr:

5−r/17=18 Step 1: Multiply both sides by 17. Step 2: Simplify both sides of the equation. Step 3: Subtract 5 from both sides. Step 4: Divide both sides by -1.